(1/2) Find the length and width of a rectangle the perimeter is 80m. The length is 4 times the width. Okay so what would the eqution be
Answer by LinnW(1048)

2W + 2L = P where W is width, L is length and P is perimeter
L = 4W Substitute 4W for L and 80 for P
2W + 2(4W) = 80
2W + 8W = 80
10W = 80
W = 8
L = 4W = 4(8) = 32